OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 5y = -10

OpenStudy (anonymous):

x-5y=-10 => 5y=x+10 => y=x/5+2 m=1/5 so slope of other line perpendicular wil be (by condition that product of slopes should be -1) -5